(5,26),(8,65),(9,82),(4,_),(7,_),(11,_) what numbers go in the blanks

Answer by Edwin McCravy(20056)

(5,26),(8,65),(9,82),(4,_),(7,_),(11,_)
 
We observe from the 1st ordered pair that 26 is 1 more than 25 which is
the square of 5.

We observe from the 2nd ordered pair that 65 is 1 more than 64 which is
the square of 8.

We observe from the 3rd ordered pair that 82 is 1 more than 81 which is
the square of 9.

So the pattern is that the second coordinate is 1 more than the square
of the first coordinate.

(4,_) The square of 4 is 16, and 16 + 1 = 17, so it's (4,17)

(7,_) The square of 7 is 49, and 49 + 1 = 50, so it's (7,50)

(11,_) The square of 11 is 121, and 121 + 1 = 122, so it's (11,122)

Edwin